I am stumped on this one, can someone help me out? I think I am overthinking this. I have literally stared at this and I have no Idea which way to go with it-- TIA

Matthew Plant has ​$36000 invested in stock A and stock B. Stock A currently sells for ​$40 a share and stock B sells for ​$60 a share. If stock B doubles in value and stock A goes up​ 50%, his stock will be worth ​$66000. How many shares of each stock does he ​own?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52775)

.
Let "n" be the number of stocks A and "m" be the number of stocks B.


Then from the condition you have this system of 2 equations in 2 unknowns


        40*n +   60*m = 36000    (1)

    1.5*40*n + 2*60*m = 66000.   (2)



Simplify


       40n +  60m = 36000        (3)

       60n + 120m = 66000        (4)


Simplify one more time by dividing all the terms by 20


       2n +  3m = 1800          (5)

       3n +  6m = 3300          (6)



Multiply eq(5) by 2 (both sides). Keep eq(6) as is.


       4n +  6m = 3600          (7)

       3n +  6m = 3300          (8)
   

Subtract eq(8) from eq(7)


        n       = 300.


Then from eq(7)   4*300 + 6m = 3600,   6m = 3600 - 1200 = 2400,   m = 2400/6 = 400.


ANSWER.  300 stocks A  and  400 stocks B.


CHECK.   You can check my solution and my answer by substituting these values m and n into the original equations.

         Do it on your own.

Solved.



Answer by MathTherapy(10551)


I am stumped on this one, can someone help me out? I think I am overthinking this. I have literally stared at this and I have no Idea which way to go with it-- TIA

Matthew Plant has ​$36000 invested in stock A and stock B. Stock A currently sells for ​$40 a share and stock B sells for ​$60 a share. If stock B doubles in value and stock A goes up​ 50%, his stock will be worth ​$66000. How many shares of each stock does he ​own?
She owns: